OpenStudy (anonymous):

Basis for the Column space of the matrix A. and Basis for the Null space. A = [1 2 1 -1] [1 2 -1 1] RREF = [1 2 1 0] [0 0 0 1] so is the column space = [1] , [-1] is this correct? [1] , [ 1] how do I get the nullspace? I know its when the matrix is multiplied with a vector = the 0 vector

OpenStudy (anonymous):

you rref is incorrect

OpenStudy (anonymous):

ok. one sec

OpenStudy (anonymous):

RREF = [1 2 0 0] [0 0 1 -1] so then the col(A) = [1] , [1] ?? [1] , [-1]

OpenStudy (anonymous):

correct, for the null space: it's [1 2 0 0 ] [x] = [0] [0 0 0 1 ] [y] = [0] yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oops correction

OpenStudy (anonymous):

would it be just [-2 1 0 0]

OpenStudy (anonymous):

null space: 1x + 1y + 0z + 0a = 0 0x + 0y + 0z + 1a = 0

OpenStudy (anonymous):

so we ge x + y = 0 and a = 0

OpenStudy (anonymous):

the matrix multiplied with the vector {-2, 1, 0 ,0} = 0 is there another vector that works?

OpenStudy (anonymous):

so x = -y [x] [-y] [-y] [0] [-1] [0] [y] = [y ] = [y] [0] y[1] z[0] [z] [z] [0] + [z] = [0] + [1] [a] [0] [0] [0] [0] [0]

OpenStudy (anonymous):

crap mistake again -.- should have been 1x + 2y + 0z + 0a = 0 0x + 0y + 1z - 1a = 0

OpenStudy (anonymous):

again, solving for x we get, x = -y and solving for z we get, z = a so (x,y,z,a) = (-y, y, a, a) so we have (-y,y,0,0) + (0,0, a,a) = y(-1,1,0,0) + a (0,0,1,1) so the basis for the null space is (-1,1,0,0) and (0,0,1,1)

OpenStudy (anonymous):

OMG! brain fart again, x = -2y? -.- I feel really stupid again.

OpenStudy (anonymous):

sall good dude I got the two vectors x2(-2,1,0,0) and x4(0,0,1,1) both = the 0 vector when multiplied with the original matrix

OpenStudy (anonymous):

yep, Don't make that stupid mistake like i did on the test. -2 is 1!!! and be sure to have your eyes check before the test XDDD

OpenStudy (anonymous):

-2 is *not* 1. That's it. I'm done for today! -.-.

OpenStudy (anonymous):

lol thanks for the help :)